# A Survey on LLM-as-a-Judge

### v4

@@ 3.1.2. Optimizing LLMs’ Output Forms
Directly requiring LLM evaluators to output evaluation results poses robustness problems. The response text may unexpectedly vary due to the inherent generative randomness of LLMs, such as outputting text like ”low relevance” while asked to measure it with discrete scores, which hinders the automated and accurate extraction of evaluation results from LLMs’ output. An effective method to enhance the robustness of output forms is to constrain LLMs’ output in structured formats within prompts. G-Eval(Liu et al., 2023a) and DHP framework(Wang et al., 2024f) perform evaluation tasks with a form-filling paradigm, constraining outputs with formats like ”X: Y”, where X represents the dimension or metric to be evaluated and Y denotes an identifiable output form like scores or specific tokens. LLM-EVAL(Lin and Chen, 2023) further modifies this form-filling paradigm, efficiently outputs evaluation results in JSON format, and obtains multidimensional scores, leveraging LLMs’ high understanding and generation capabilities of code-like textural formats.
@@ 3.2. Improvement Strategy of LLMs’ Abilities
The evaluation capabilities of LLMs are a reflection of their powerful general language understanding and generation abilities triggered by specific prompts. Methods for optimizing evaluation through prompt design, which focuses on LLMs’ in-contextual learning capabilities, require LLMs to fully comprehend the meaning of prompts and consistently follow the relevant evaluation instructions. However, even state-of-the-art LLMs like GPT-4 face issues such as conceptual confusion(Hu et al., 2024), and smaller open-source LLMs have even more limitations in their evaluation capabilities. Therefore, how to improve the evaluation capabilities of LLMs, including how to fine-tune LLMs through meta-evaluation datasets and how to iteratively optimize models based on the feedback of evaluation results, is significant for improving the fundamental evaluation performance of LLM-as-a-judge.
@@ 3.2.1. Fine-tuning via Meta Evaluation Datasets
A straightforward approach to enhancing the evaluation capabilities of LLMs is to fine-tune them via meta-evaluation datasets specifically constructed for evaluation tasks, which helps improve the LLMs’ understanding of specific evaluation prompts, boosts the evaluation performance, or addresses potential biases. The most critical step in this optimization strategy is the collection and construction of training data. A common method involves sampling evaluation questions from publicly available datasets, modifying them with certain templates, and supplementing the dataset with evaluation responses generated either manually or by powerful LLMs like GPT4. For instance, PandaLM(Wang et al., 2023d) samples inputs and instructions from Alpaca 52K(Taori et al., 2023) and generate responses using GPT-3.5 to construct training data, while SALAD-Bench(Li et al., 2024c) builds its training data from a subset of LMSYS-Chat(Zheng et al., 2023a) and Toxicchat(Lin et al., 2023).
@@ 3.2.2. Iterative Optimization Based on Feedback of Evaluation Results
A typical example is INSTRUCTSCORE(Xu et al., 2023). To improve model performance and further benefit the final quality score calculation, this scoring framework collects failure modes of metric outputs, queries GPT-4 on each failure mode to gather automatic feedback, and finally selects explanations most aligned with human preferences to iteratively fine-tune the LLaMA model. Unlike INSTRUCTSCORE which directly optimizes the model, the LLM evaluator in JADE(Zhang et al., 2023b) relies on human judges to correct LLMs’ evaluation results and updates the most frequently corrected samples into the example sets for few-shot prompting. JADE utilizes this relatively low-cost method to achieve iterative updates of the evaluation capabilities.
@@ 3.3.1. Integration of Multiple Evaluation Results
In addition to integrating results from multiple rounds of evaluation, using multiple LLM evaluators to assess the contents simultaneously and integrating the results is another effective method, which can reduce biases introduced by LLMs. For instance, CPAD(Liu et al., 2023d) utilizes ChatGLM-6B(Du et al., 2022), Ziya-13B(Zhang et al., 2022), and ChatYuan-Large-v2(Zhang and Zhao, 2022) as evaluators to evaluate the contents and obtain the final results by voting. Bai et al.(Bai et al., 2023b) propose a novel evaluation method called decentralized peer review of LLMs, which utilizes LLMs that generate contents to evaluate each other’s generated contents and eventually integrate the results.
@@ 4. Evaluation of LLM Evaluators
Despite their impressive performance, LLMs exhibit several notable shortcomings, such as hallucinations (Tonmoy et al., 2024), biases (Gallegos et al., 2024a), and a lack of robustness (Zhu et al., 2023d). When LLMs are employed as evaluators, these inherent issues can lead to suboptimal evaluation outcomes. Therefore, it is crucial to accurately and comprehensively assess the quality of LLM-as-a-judge and identify potential vulnerabilities. This section will review existing work on the evaluation of LLM-as-a-judge, focusing on three key areas: base metric (Section 4.1), bias (Section 4.2), and robustness (Section 4.3).
@@ 4.1. Basic Metric
where $\mathcal{D}$ is the dataset, $S_{\text{llm}}$ and $S_{\text{human}}$ is the evaluation result of LLM evaluator and human judge respectively, which can be in the form of both score or rank. Additionally, widely used correlation metrics such as Cohen’s Kappa (Thakur et al., 2024) and Spearman’s correlation (Liu et al., 2024b; Bai et al., 2023b) are also employed to access agreement. Other works treat the LLM-as-a-judge task as a classification problem, where human annotations serve as the labels, and compute precision, recall, and F1 scores to evaluate the performance (Zhu et al., 2023b; Wang et al., 2023d).
Both of above metrics rely on the datasets with LLM-generated response and responding human judgements. Therefore, there is also a practical need to construct a comprehensive benchmark for the meta-evaluation. In (Zheng et al., 2023d), MTBench and Chatbot Arena Conversations are proposed. The former has only 80 human-crafted queries, each with several LLMs’ responses and expert-level human annotation on pairwise comparison; the latter is a large collection of crowdsourced data, with more than 30K queries from real-world users and their vote on pairs of responses from different LLMs. FairEval (Wang et al., 2023b) is based on the 80 queries from VicunaBench (Touvron et al., 2023b) with human annotated labels between ChatGPT and Vicuna responses. PandaLM (Wang et al., 2023d) constructs a test set comprising 999 pairwise samples, with queries from 252 user-oriented instructions in (Wang et al., 2022). LLMEval2 (Zhang et al., 2023c) is much larger than the previous two, with 2,553 samples compiled from multiple data sources with human-annotated preferences. Shepherd (Wang et al., 2023c) collects 352 samples from multiple sources for its critique model as a test set to evaluate the quality of the critiques. In (Zeng et al., 2023), a meta-evaluation benchmark was created to evaluate the effectiveness of LLM evaluators in recognizing instruction-following outputs, focusing on only a part of the LLM-as-a-Judge pipeline.
More recent works propose benchmark for code (Zhao et al., 2024b), multi-modal (Chen et al., 2024b) and non-English (Son et al., 2024) tasks. Table 1 shows the benchmarks and their statistics.
@@
Current meta-evaluation primarily focuses on LLM-as-a-judge for models, while there is a lack of sufficient meta-evaluation when these LLM evaluators are used for automatically annotating large-scale datasets (Section 2.4.2). We advocate for more rigorous accessment of the alignment between LLM-as-a-judge and human judgment when they are employed for large-scale data annotation. Additionally, it is also crucial to assess the potential bias and robustness, which will be discussed in the following sections.
@@
Benchmark Release Year Size Annotation Format Evaluation Dimension Agreement Position Bias Length Bias Bias Types MTBench (Zheng et al., 2023d) 2023 80 Pairwise 3 Chatbot Arena (Zheng et al., 2023d) 2023 30k Pairwise 3 FairEval (Wang et al., 2023b) 2023 80 Pairwise 1 PandaLM (Wang et al., 2023d) 2023 - Pairwise 0 LLMEval2 (Zhang et al., 2023c) 2023 2553 Pairwise 0 Shepherd (Wang et al., 2023c) 2023 1317 Score 0 EvalBiasBench (Park et al., 2024) 2023 80 Pairwise 6 CALM (Ye et al., 2024) 2024 4356 Pairwise & Score 12 JudgeBench (Tan et al., 2024b) 2024 - Pairwise 0 MLLM-as-a-Judge (Chen et al., 2024b) 2024 30k Pairwise & Score 0 CodeJudge (Zhao et al., 2024b) 2024 1860 Score 0 KUDGE (Son et al., 2024) 2024 3324 Pairwise & Score 0
@@ 4.2. Bias
Previous reviews have highlighted that large language models exhibit various types of biases across various tasks (Gallegos et al., 2024b; Dai et al., 2024; Tan et al., 2024a). These internal biases of LLMs may also affect LLM-as-a-judge, leading to unfair evaluation outcomes and subsequently impacting the development of LLMs. Therefore, it is crucial to understand the types of biases that LLM evaluators might possess and to systematically assess these biases. In this section, we systematically review various types of biases in the LLM-as-a-judge context, including their definitions, relevant metrics, and datasets that can be used for evaluation.
@@
Length Bias refers to the tendency to favor responses of a particular length, such as a preference for more verbose responses which is also known as verbosity bias (Park et al., 2024; Huang et al., 2024; Zheng et al., 2023d; Ye et al., 2024). Length bias can be revealed by rephrasing one of the original response into a more verbose one (Ye et al., 2024; Zheng et al., 2023d). Even though these expansions do not introduce new information, there is still concern regarding changes to the original response in terms of perplexity, fluency, or style. Alternatively, previous study (Saito et al., 2023) investigated this bias by comparing multiple sampled responses and revealed a statistical tendency towards longer answers. However, ensuring the comparable quality of multiple samples remains a challenging problem.
Self-Enhancement Bias describe the phenomenon that LLM evaluators may prefer response generated by themselves (Zheng et al., 2023d; Ye et al., 2024). Considering the significant self-enhancement bias, as suggested in (Ye et al., 2024), we should avoid using the same model as the evaluator. This is only a stopgap, as we may not use the optimal evaluator when evaluating the most advanced LLMs.
Other Bias. Diversity Bias refers to bias against certain demographic groups (Ye et al., 2024), including certain genders (Chen et al., 2024a), race, and sexual orientation (Kumar et al., 2024b). As revealed in (Chen et al., 2024a), evaluators’ tendency toward visually appealing content, regardless of its actual validity, such as the text with emoji. Concreteness bias reflects that LLM evaluators favor responses with specific details, including citation of authoritative sources, numerical values and complex terminologies, which is called authority bias (Park et al., 2024) or citation bias (Chen et al., 2024a; Ye et al., 2024). Furthermore, LLM evaluators may favor response with certain emotional tones, such as cheerful, sad, angry, and fearful, which is defined as seyiment bias (Ye et al., 2024; Li and Sinnamon, 2023).

### v5

@@ 3.1.2. Optimizing LLMs’ Output Forms
Directly requiring LLM evaluators to output evaluation results poses robustness problems. The response text may unexpectedly vary due to the inherent generative randomness of LLMs, such as outputting text like ”low relevance” while asked to measure it with discrete scores, which hinders the automated and accurate extraction of evaluation results from LLMs’ output. An effective method to enhance the robustness of output forms is to constrain LLMs’ output in structured formats within prompts. G-Eval(Liu et al., 2023b) and DHP framework(Wang et al., 2024f) perform evaluation tasks with a form-filling paradigm, constraining outputs with formats like ”X: Y”, where X represents the dimension or metric to be evaluated and Y denotes an identifiable output form like scores or specific tokens. LLM-EVAL(Lin and Chen, 2023) further modifies this form-filling paradigm, efficiently outputs evaluation results in JSON format, and obtains multidimensional scores, leveraging LLMs’ high understanding and generation capabilities of code-like textural formats.
@@ 3.2. Improvement Strategy of LLMs’ Abilities
The evaluation capabilities of LLMs are a reflection of their powerful general language understanding and generation abilities triggered by specific prompts. Methods for optimizing evaluation through prompt design–focused on LLMs’ in-contextual learning capabilities–require LLMs to fully comprehend the meaning of prompts and consistently follow the relevant evaluation instructions. However, even state-of-the-art LLMs like GPT-4 encounter problems such as conceptual confusion(Hu et al., 2024), and smaller open-source LLMs have even more limitations in their evaluation capabilities. Consequently, refining the evaluation capabilities of LLMs, including how to fine-tune LLMs through meta-evaluation datasets and how to iteratively optimize models based on the feedback of evaluation results, is significant for improving the fundamental evaluation performance of LLM-as-a-judge.
@@ 3.2.1. Fine-tuning via Meta Evaluation Datasets
A straightforward approach to enhancing the evaluation capabilities of LLMs is to fine-tune them via meta-evaluation datasets specifically constructed for evaluation tasks, which helps improve the LLMs’ understanding of specific evaluation prompts, boosts the evaluation performance, or addresses potential biases. The most critical step in this optimization strategy is the collection and construction of training data. A common method involves sampling evaluation questions from publicly available datasets, modifying them with certain templates, and supplementing the dataset with evaluation responses generated either manually or by powerful LLMs like GPT4. For instance, PandaLM(Wang et al., 2023c) samples inputs and instructions from Alpaca 52K(Taori et al., 2023) and generates responses using GPT-3.5 to construct training data, while SALAD-Bench(Li et al., 2024b) builds its training data from a subset of LMSYS-Chat(Zheng et al., 2023a) and Toxicchat(Lin et al., 2023).
@@ 3.2.2. Iterative Optimization Based on Feedback of Evaluation Results
A typical example is INSTRUCTSCORE(Xu et al., 2023b). To improve model performance and further benefit the final quality score calculation, this scoring framework collects failure modes of metric outputs, queries GPT-4 on each failure mode to gather automatic feedback, and finally selects explanations most aligned with human preferences to iteratively fine-tune the LLaMA model. Unlike INSTRUCTSCORE which directly optimizes the model, the LLM evaluator in JADE(Zhang et al., 2023c) relies on human judges to correct LLMs’ evaluation results and updates the most frequently corrected samples into the example sets for few-shot prompting. JADE utilizes this relatively low-cost method to achieve iterative updates of the evaluation capabilities.
@@ 3.3.1. Integration of Multiple Evaluation Results
In addition to integrating results from multiple rounds of evaluation, using multiple LLM evaluators to assess the contents simultaneously and integrating the results is another effective method, which can reduce biases introduced by LLMs. For instance, CPAD(Liu et al., 2023e) utilizes ChatGLM-6B(Du et al., 2022), Ziya-13B(Zhang et al., 2022), and ChatYuan-Large-v2(Zhang and Zhao, 2022) as evaluators to evaluate the contents and obtain the final results by voting. Bai et al.(Bai et al., 2023) propose a novel evaluation method called decentralized peer review of LLMs, which utilizes LLMs that generate contents to evaluate each other’s generated contents and eventually integrate the results.
@@ 4. Evaluation of LLM Evaluators
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw, rounded corners, align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=1pt, , where level=1text width=5em,fill=blue!10, where level=2text width=5em,font=,fill=pink!30, where level=3font=,yshift=0.26pt,fill=yellow!20, [Meta
-Evaluation, fill=green!20 [Base Metric
(Sec. 4.1), text width=8em [Agreement (Thakur et al., 2024) Cohen’s Kappa (Thakur et al., 2024)
Spearman’s correlation  (Liu et al., 2024b; Bai et al., 2023), text width=15em ] ] [Bias (Sec. 4.2), text width=8em [Task-Agnostic Biases
(Sec. 4.2.1), text width=7em [Diversity Bias (Ye et al., 2024) Cultural Bias
Self-Enhancement Bias (Zheng et al., 2023b; Ye et al., 2024)] ] [Judgment-Specific Biases
(Sec. 4.2.2), text width=8em [Position Bias(Wang et al., 2024c; Thakur et al., 2024; Shi et al., 2024; Ye et al., 2024)
Compassion-fade bias (Koo et al., 2023; Ye et al., 2024)
Style Bias (Chen et al., 2024a; Ye et al., 2024; Li and Sinnamon, 2023)
Length Bias (Park et al., 2024; Huang et al., 2024; Zheng et al., 2023b; Ye et al., 2024)
Concreteness Bias (Park et al., 2024; Chen et al., 2024a; Ye et al., 2024)] ] ] [Adversarial Robustness
(Sec. 4.3), text width=9em [Adversarial Phrases Attack (Raina et al., 2024), text width=11em ] [Null Model Attack (Zheng et al., 2024a), text width=8em ] [Majority Opinions Attack (Ye et al., 2024; Koo et al., 2023), text width=13em ] [Meaningless Statement Robustness (Ye et al., 2024; Koo et al., 2023), text width=15em ] ] ]
Figure 13. Structure of Meta-Evaluation.
Despite their impressive performance, LLMs exhibit several notable shortcomings, such as hallucinations (Tonmoy et al., 2024), biases (Gallegos et al., 2024a), and a lack of robustness (Zhu et al., 2023c). When LLMs are employed as evaluators, these inherent issues can lead to suboptimal evaluation outcomes. Therefore, it is crucial to accurately and comprehensively assess the quality of LLM-as-a-judge and identify potential vulnerabilities. This section will review existing work on the evaluation of LLM-as-a-judge, focusing on three key areas: base metric (Section 4.1), bias (Section 4.2), and robustness (Section 4.3).
@@ 4.1. Basic Metric
where $\mathcal{D}$ is the dataset, $S_{\text{llm}}$ and $S_{\text{human}}$ is the evaluation result of LLM evaluator and human judge respectively, which can be in the form of both score or rank. Additionally, widely used correlation metrics such as Cohen’s Kappa (Thakur et al., 2024) and Spearman’s correlation (Liu et al., 2024b; Bai et al., 2023) are also employed to access agreement. Other works treat the LLM-as-a-judge task as a classification problem, where human annotations serve as the labels, and compute precision, recall, and F1 scores to evaluate the performance (Zhu et al., 2023a; Wang et al., 2023c).
Datasets. Both of above metrics rely on the datasets with LLM-generated response and responding human judgments. Therefore, there is also a practical need to construct a comprehensive benchmark for the meta-evaluation. We list the existing benchmarks and their statistics in Table 1. MTBench (Zheng et al., 2023b) has only 80 human-crafted queries with their responding human annotation and LLMs’ responses. FairEval (Wang et al., 2024c) is constructed from the 80 queries from VicunaBench (Touvron et al., 2023b) with human annotated preference between ChatGPT and Vicuna responses. Chatbot Arena Conversations (Zheng et al., 2023b) is a larger collection of crowdsourced data (about 30k) with human annotated preference. Research (Zeng et al., 2023) construct a benchmark to access the capability of LLM evaluator in evaluating whether a response is following the instruction. This dataset contains human-curated 419 pairs of outputs, one adhering to instructions while the other diverging, yet may possess deceptive qualities that mislead an LLM evaluator. Research (Chen et al., 2024b) evaluate the capabilities of multi-modal LLMs in assisting evaluation tasks across various modalities and introduce MLLM-as-a-Judge, a comprehensive multi-modal benchmark. Recent advances also expand the scope of meta-evaluation benchmarks to specialized domains, including code assessment (Zhao et al., 2024b) and non-English language tasks (Son et al., 2024). Furthermore, CALM (Ye et al., 2024) presents a systematic framework for bias quantification, featuring an automated perturbation mechanism to generate meta-evaluation data for examining 12 distinct types of potential biases in LLM evaluators.
@@
Current meta-evaluation primarily focuses on LLM-as-a-judge for models, while there is a lack of sufficient meta-evaluation when these LLM evaluators are used for automatically annotating large-scale datasets (Section 2.4.2). We advocate for more rigorous assessment of the alignment between LLM-as-a-judge and human judgment when they are employed for large-scale data annotation. Additionally, it is also crucial to assess the potential bias and robustness, which will be discussed in the following sections.
@@
Benchmark Release Year Size Annotation Format Evaluation Dimension Agreement Position Bias Length Bias Bias Types MTBench (Zheng et al., 2023b) 2023 80 Pairwise 3 Chatbot Arena (Zheng et al., 2023b) 2023 30k Pairwise 3 FairEval (Wang et al., 2024c) 2023 80 Pairwise 1 PandaLM (Wang et al., 2023c) 2023 - Pairwise 0 LLMEval2 (Zhang et al., 2023d) 2023 2553 Pairwise 0 Shepherd (Wang et al., 2023b) 2023 1317 Score 0 EvalBiasBench (Park et al., 2024) 2023 80 Pairwise 6 CALM (Ye et al., 2024) 2024 4356 Pairwise & Score 12 JudgeBench (Tan et al., 2024b) 2024 - Pairwise 0 MLLM-as-a-Judge (Chen et al., 2024b) 2024 30k Pairwise & Score 0 CodeJudge (Zhao et al., 2024b) 2024 1860 Score 0 KUDGE (Son et al., 2024) 2024 3324 Pairwise & Score 0
@@ 4.2. Bias
Previous reviews have highlighted that large language models exhibit various types of biases across various tasks (Gallegos et al., 2024a; Dai et al., 2024a; Tan et al., 2024a). These internal biases of LLMs may also affect LLM-as-a-judge, leading to unfair evaluation outcomes and subsequently impacting the development of LLMs. Therefore, it is crucial to understand the types of biases that LLM evaluators might possess and to systematically assess these biases. In this section, we systematically review various types of biases in the LLM-as-a-judge context, including their definitions, relevant metrics, and datasets that can be used for evaluation.
The meta-evaluation of LLM-as-a-judge introduces systematic biases that can be broadly categorized into two classes: task-agnostic biases inherent to LLMs across general applications, and judgment-specific biases unique to LLM-as-a-judge scenarios. This taxonomy aims to clarify their distinct characteristics and implications.
4.2.1. Task-Agnostic Biases
These biases manifest across diverse LLM applications, including open-domain QA, classification, and summarization. However, when arising in the LLM-as-a-judge, the biases are particularly critical due to their cascading effects on downstream tasks. When LLM-generated judgments serve as feedback for model training or data annotation, these biases risk being amplified and propagated. We present a few typical examples and recommend consulting comprehensive reviews on language model bias (Gallegos et al., 2024b; Guo et al., 2024) for a more thorough understanding.
Diversity Bias refers to bias against certain demographic groups (Ye et al., 2024), including certain genders (Chen et al., 2024a), race, and sexual orientation (Kumar et al., 2024b). In the context of LLM-as-a-judge scenarios, this bias may appear when evaluators give higher scores to responses that align with stereotypes of certain groups.
Cultural Bias. In general domains, cultural bias refers to situations where models might misinterpret expressions from different cultures or fail to recognize regional language variants (Gallegos et al., 2024b). In the context of LLM-as-a-judge, it indicates that evaluators might score expressions from unfamiliar cultures poorly.
Self-Enhancement Bias describe the phenomenon that LLM evaluators may prefer response generated by themselves (Zheng et al., 2023b; Ye et al., 2024). This bias has also been known as source bias in retrieval task (Dai et al., 2024b) and open-domain question answering systems (Tan et al., 2024a). Considering the significant self-enhancement bias, as suggested in (Ye et al., 2024), we should avoid using the same model as the evaluator. This is only a stopgap, as we may not use the optimal evaluator when evaluating the most advanced LLMs.
4.2.2. Judgment-Specific Biases
Judgment-specific biases are either unique to the LLM-as-a-judge setting or have a significant impact on judgment tasks. A classic example is the ”position bias”, which has a more pronounced effect in the context of LLM-as-a-judge where the evaluator often need to compare pairwise responses. Different from task-agnostic biases, judgment-specific biases are more difficult to resolve naturally with the development of foundational large model capabilities and require targeted optimization for judgment tasks.
@@ 4.2.2. Judgment-Specific Biases
Compassion-fade bias describes the effect of the model names (Koo et al., 2023; Ye et al., 2024). This tendency occurs when we explicitly provide model names; for instance, evaluators may be inclined to give higher scores to results labeled as “gpt-4”. This tendency underscores the necessity of anonymous evaluation.
Style Bias refers to the tendency towards certain text style. As revealed in (Chen et al., 2024a), evaluator may also prefer visually appealing content, regardless of its actual validity, such as the text with emoji. Furthermore, LLM evaluators may favor response with certain emotional tones, such as cheerful, sad, angry, and fearful, which is defined as sentiment bias (Ye et al., 2024; Li and Sinnamon, 2023).
Length Bias refers to the tendency to favor responses of a particular length, such as a preference for more verbose responses which is also known as verbosity bias (Park et al., 2024; Huang et al., 2024; Zheng et al., 2023b; Ye et al., 2024). Length bias can be revealed by rephrasing one of the original response into a more verbose one (Ye et al., 2024; Zheng et al., 2023b). Even though these expansions do not introduce new information, there is still concern regarding changes to the original response in terms of perplexity, fluency, or style. Alternatively, previous study (Saito et al., 2023) investigated this bias by comparing multiple sampled responses and revealed a statistical tendency towards longer answers. However, ensuring the comparable quality of multiple samples remains a challenging problem.
Concreteness Bias reflects that LLM evaluators favor responses with specific details, including citation of authoritative sources, numerical values and complex terminologies, which is called authority bias (Park et al., 2024) or citation bias (Chen et al., 2024a; Ye et al., 2024).
The negative effects of concreteness bias arises from the neglect of the factual correctness of these details, thereby encouraging hallucination (Agrawal et al., 2023).
4.2.3. Challenges